OpenStudy (anonymous):

find the 2nd derivitave of sinx/3+cos2x but please show me the first step !!

OpenStudy (anonymous):

\[\frac{\sin(x)}{3+\cos(2x)} ?\]

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

whats my first step?? like im having trouble with the 3

OpenStudy (anonymous):

my first step i got is cosx (3+cos2x) - sin(-sin2x)(s)/ (3+cos2x)^2

OpenStudy (anonymous):

The following is the first derivative I applied the quotient rule because we had a quotient \[\frac{(\sin(x))' \cdot (3+\cos(2x)-(\sin(x)) \cdot (3+\cos(2x))'}{(3+\cos(2x))^2}\]

OpenStudy (anonymous):

\[\frac{\cos(x) \cdot (3+\cos(2x))-\sin(x) \cdot (0-2\sin(2x))}{(3+\cos(2x))^2}\]

OpenStudy (anonymous):

so i was right?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

some of what you wrote is right

OpenStudy (anonymous):

ok, ill just ask my teacher tomo, cause this is a long question :p

OpenStudy (anonymous):

aww...you don't understand what I did?

OpenStudy (anonymous):

well, i did, but i need to find 2nd derivative and thats gonna take a loooong time :p

OpenStudy (anonymous):

\[\frac{\cos(x)(3+\cos(2x))+2\sin(2x)\sin(x)}{(3+\cos(2x))^2}\] \[\frac{3\cos(x)+\cos(2x)\cos(x)+2\sin(2x)\sin(x)}{(3+\cos(2x))^2}\] ------------- Now second derivative Again we need to do quotient rule

OpenStudy (anonymous):

.............. how did u do the 2nd part???

OpenStudy (anonymous):

lol It is too long to fit on one line one sec

OpenStudy (anonymous):

Gosh I can't make it look pretty on here. You have to apply quotient rule first

OpenStudy (anonymous):

Then you will have to use product rule a couple of times

OpenStudy (anonymous):

and chain rule too

OpenStudy (anonymous):

How about this we do it in small parts

OpenStudy (anonymous):

its okay man its alot of work :) ill do it tomo with my teacher, dont waste your time

OpenStudy (anonymous):

\[(\frac{3\cos(x)+\cos(2x)\cos(x)+2\sin(2x)\sin(x)}{(3+\cos(2x))^2})'\] \[=\frac{(3\cos(x)+\cos(2x)\cos(x)+2\sin(2x)\sin(x))'(3+\cos(2x))^2}{(3+\cos(2x))^4} \] \[-\frac{(3\cos(x)+\cos(2x)\cos(x)+2\sin(2x)\sin(x))((3+\cos(2x))^2)'}{(3+\cos(2x))^4}\]

OpenStudy (anonymous):

Couldn't put it all on one line So we need to find this: \[(3\cos(x)+\cos(2x)\cos(x)+2\sin(2x)\sin(x))'\] and \[((3+\cos(2x))^2)'\]

OpenStudy (anonymous):

\[(3\cos(x)+\cos(2x)\cos(x)+2\sin(2x)\sin(x))'\] \[=-3\sin(x)+(-2\sin(2x)\cos(x)+\cos(2x)(-\sin(x))+2(2\cos(2x)\sin(x) \] \[+\sin(2x)(\cos(x))\]

OpenStudy (anonymous):

\[((3+\cos(2x))^2)'=2(3+\cos(2x))(0-2\sin(2x))\] Now just plug in these parts :)

OpenStudy (anonymous):

You might want to simplify too if possible